OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ Numerical Answers Expected!

OpenStudy (kirbykirby):

Try plugging in 5 into f(n): \[f(5)=f(1)+f(2)+f(5−1)=2+4+f(4)=6+f(4)\] what is f(4)? use the rule again \[ f(4)=f(1)+f(2)+f(4−1)=2+4+f(3)=6+f(3)\] And then \[f(3)=f(1)+f(2)+f(3−1)=2+4+f(2)=6+4=10\] So you now know f(3)'s value. You can go back to f(4) and find its value, and then go back to f(5) and find its value since you'll know f(4)